# Hierarchical Learning and Computing over Space-Ground Integrated Networks

### v1

@@ IV-B4 Impact of LEO Constellations
First of all, with the same number of LEO satellites, the system with Walker-Delta constellation exhibits lower average energy consumption compared to the Walker-Star constellation due to the closer inter-satellite distances at lower orbital altitudes, resulting in stronger received signal strength. Secondly, as the number of LEO satellites increases, both the density of satellites within the same orbit and the density of orbital distribution increase. This results in significantly reduced inter-satellite distances within and between orbits, substantially enhancing ISL receiving power. Consequently, the averaged transmission energy required to complete a single global model aggregation is significantly reduced, and the outage probability is also notably decreased. This demonstrates the effectiveness of network scaling. Besides, from Table II, we observe that the proposed TAEER outperforms the D-Merge method regardless of whether the LEO satellite network is dense or sparse. Notably, the advantage of TAEER is even more pronounced in sparse networks, highlighting the robustness of the proposed TAEER.
@@ V Conclusion
In this paper, we introduced a novel hierarchical learning and computing framework tailored for space-ground integrated networks, which effectively supports collaborative training on widely distributed terrestrial IoT devices using LEO mega-constellations and GEO satellites. We leveraged the predictability of satellite network topology by modeling the space network as a directed graph, where edge weights represent energy consumption during transmission. We revealed that minimizing the overall network transmission energy is equivalent to solving a DST problem, which is NP-hard. To address this, we proposed a TAEER algorithm. By first using the Dijkstra algorithm to find the minimum cost path to the root node and then identifying a MSA, we solve the DST problem heuristically to create an effective aggregation routing path. Furthermore, we extended our solution to account for unpredictability such as transmission outages. Simulation results confirmed its superior performance and robustness compared to existing benchmarks.
Future Directions: There are several open directions for extending this work. For instance, this paper considers the space topology as quasi-static, meaning it varies between different time slots but remains constant within the same time slot, while the connection weights are time-varying within a time slot. A promising research direction is to model the space network as a time-varying graph for analysis [38]. Additionally, we can explore the use of deep reinforcement learning and other low-complexity methods to solve the DST problem for topology-aware routing.

### v2

@@ IV-B4 Impact of LEO Constellations
First of all, with the same number of LEO satellites, the system with Walker-Delta constellation exhibits lower average energy consumption and outage probability compared to the Walker-Star constellation due to the closer inter-satellite distances at lower orbital altitudes, resulting in stronger received signal strength. Secondly, as the number of LEO satellites increases, both the density of satellites within the same orbit and the density of orbital distribution increase. This results in significantly reduced inter-satellite distances within and between orbits, substantially enhancing ISL receiving power. Consequently, the averaged transmission energy required to complete a single global model aggregation is significantly reduced from $80/4/1$ to $240/8/1$ constellation. We can also observe that as the number of satellites increases further, the averaged transmission energy of the $800/20/1$ constellation is slightly higher compared to the $240/8/1$ constellation due to the more frequent relay switching of ISL hops. Besides, the outage probability is notably decreased as the total number of satellites increases, which demonstrates the effectiveness of network scaling. Besides, from Table II, we observe that the proposed TAEER outperforms the D-Merge method regardless of whether the LEO satellite network is dense or sparse. Notably, the advantage of TAEER is even more pronounced in sparse networks, highlighting the robustness of the proposed TAEER.
@@ V Conclusion
In this paper, we introduced a novel hierarchical learning and computing framework tailored for SGINs, which effectively supports collaborative training on widely distributed terrestrial IoT devices using LEO mega-constellations and GEO satellites. We leveraged the predictability of satellite network topology by modeling the space network as a directed graph, where edge weights represent energy consumption during transmission. We revealed that minimizing the overall network transmission energy is equivalent to solving a DST problem, which is NP-hard. To address this, we proposed a TAEER algorithm. By first using the Dijkstra algorithm to find the minimum cost path to the root node and then identifying a MSA, we solve the DST problem heuristically to create an effective aggregation routing path. Furthermore, we extended our solution to account for unpredictability such as transmission outages. Simulation results confirmed its superior performance and robustness compared to existing benchmarks.
Future Directions: There are several open directions for extending this work. First of all, detailed modeling of the energy harvesting and battery systems of the satellites can be considered to formulate a constrained DST problem. Second, this paper considers the space topology as quasi-static, meaning it varies between different time slots but remains constant within the same time slot, while the connection weights are time-varying within a time slot. A promising research direction is to model the space network as a time-varying graph for analysis [32]. We also plan to explore allocating the proportion of transmitted information across time frames, which represents a data allocation problem under dynamic topology. Additionally, leveraging the correlation between topologies in adjacent time slots to reduce the complexity of topology construction is another important potential research direction. Finally, we can explore the use of deep reinforcement learning and other low-complexity methods to solve the DST problem for topology-aware routing.